# Efficient Robot Design with Multi-Objective Black-Box Optimization and Large Language Models

### v1

@@ I INTRODUCTION
As illustrated, various methods such as genetic algorithms, evolutionary strategies, exhaustive search, heuristic search, and GANs have been utilized for optimizing the body design of diverse robots. Among these, research using black-box optimization, particularly genetic algorithms, is prevalent due to its applicability to complex problem settings, including optimizations involving both continuous and discrete variables as well as multi-objective optimization [yang2000modular, xiao2016designopt, kawaharazuka2023autodesign, kawaharazuka2024slideopt, vanteddu2024cadurdf]. Although gradient-free black-box optimization is applicable to any problem, it has the drawback of requiring a substantial number of samples and time to generate high-quality solutions.
@@
To address this, we propose a method that combines black-box optimization with LLMs to enable more efficient output of design solutions. Our approach leverages the strength of black-box optimization for precise numerical adjustments and the capability of LLMs to produce broad design outputs. In parallel with the sampling process in conventional black-box optimization, sampling is also performed using an LLM, which is provided with prior knowledge, problem settings, and design examples. It should be noted that while the combination of LLMs and BBO has been shown to be effective for single-objective and general numerical problems [liu2024large], this study extends the approach to the more complex domain of multi-objective optimization in robot design. In this study, a multi-objective design optimization is conducted by adjusting the robot’s base position, joint configuration, and link lengths to minimize the end-effector position error and the required torque for given operation points. We demonstrate that our method enables more efficient exploration of design solutions compared to black-box optimization alone, while also discussing the method’s interesting characteristics and limitations.
@@ II Efficient Robot Design with Multi-Objective Black-Box Optimization and LLM-based Sampling
In Section II-A, we describe the design parameters of the robot used in this study. In Section II-B, we explain the objective functions to be optimized. In Section II-C, we provide details on the sampling process using multi-objective black-box optimization. In Section II-D, we describe the sampling process that leverages LLMs.
@@ II-B Objective Functions
where $\bm{P}^{ik}_{i}$ and $\bm{\tau}^{ik}_{i}$ represent the end-effector position and gravity-compensation torque, respectively, when the robot solves the inverse kinematics (IK) to reach the target position $\bm{P}^{ref}_{i}$, assuming the robot’s initial posture has all joint angles set to zero. The inverse kinematics algorithm used in this study is based on the methods described in [chan1995ik, sugiura2007ik]. In other words, this evaluates whether the target end-effector position can be achieved with that body design, and how much joint torque is required in doing so. Of course, it is also possible to consider not only the target position but also the orientation, or to use a more complex evaluation function.
@@ II-C Multi-Objective Black-Box Optimization
In this study, we address a multi-objective optimization problem that aims to minimize both objective functions $E^{pos}$ and $E^{torque}$. This allows us to obtain various design solutions from the Pareto-optimal set. For the multi-objective optimization, we use the Tree-Structured Parzen Estimator (TPE) [bergstra2011tpe] implemented in the black-box optimization library Optuna [akiba2019optuna]. TPE supports multi-objective optimization and is notable for requiring a relatively small recommended sample size of fewer than 1000. Although NSGA-II [deb2002nsgaii] is also commonly used, this study adopted TPE, which can obtain good solutions with fewer samples.
@@ III-A Experimental Setup
In this study, experiments are conducted for three sets of target operation points, Target-1, Target-2, and Target-3, as shown in Fig. 6. For each target operation point, we compare BBO, which uses only black-box optimization, and BBO/LLM+, which combines black-box optimization with a large language model. To evaluate the extent of LLM’s contribution to the optimization process, we vary $N_{step}$, which represents the number of black-box optimization steps performed between each LLM-based sampling. In this study, experiments are conducted with $N_{step}={5,10,50}$. Additionally, we perform a comparative evaluation of BBO/LLM-, which uses general Chain-of-Thought reasoning without specific parameter analysis instructions. The other parameters are set as follows: $D=4$, $N_{init}=10$, $N_{pareto}=5$, $N_{random}=5$, and $N_{total}=200$. Each experiment is repeated five times with different random seed values, and the mean and standard deviation of the results are reported.
@@ III-B Experimental Results
The experimental results are presented below. Fig. 7 shows a comparison of the methods BBO, BBO/LLM+ ($N_{step}=5$), BBO/LLM+ ($N_{step}=10$), and BBO/LLM+ ($N_{step}=50$) for Target-1, Target-2, and Target-3. The horizontal axis represents the optimization iterations, and the vertical axis represents the hypervolume $E_{hv}$. The figure illustrates how $E_{hv}$ changes with different values of $N_{step}$. A general trend observed is that BBO/LLM+ ($N_{step}=10$) consistently demonstrates the best results across all target operation points, confirming the effectiveness of LLM-based sampling. For Target-1, the improvement is particularly significant, while for Target-2 and Target-3, the impact is less pronounced, indicating that the effectiveness of LLM sampling depends on the target operation points. For $N_{step}=50$, only four LLM-based samplings occur within 200 sampling steps, making the trend of $E_{hv}$ similar to that of BBO. For $N_{step}=5$, although LLM-based sampling occurs more frequently and is effective in the early stages for Target-1 and Target-2, the results are inferior compared to $N_{step}=10$.
@@ III-D Target-2 Experiment
We describe the experimental results for Target-2 (Fig. 11). In the case of Target-2, LLM-based sampling also accounts for many of the Pareto-optimal solutions, confirming its effectiveness. Unlike Target-1, the joint structure for both Solution 1 and Solution 2 is the same, YRPRY-R-P-R. However, the second and fourth links in Solution 1 are longer compared to those in Solution 2. Due to this difference, Solution 1 achieves a lower $E^{pos}$, while Solution 2 achieves a lower $E^{torque}$. In Solution 2, the lengths of the first and second links are set to approximately 0.5 m, allowing the second link to match the height of the target operation points. The remaining joints are then used in a manner similar to a SCARA-type structure, helping to minimize the required torque.
@@ IV Discussion
Finally, we discuss the limitation of the proposed method. In this study, we adopted a relatively straightforward and general problem setting. However, in real-world robot design, it is often necessary to consider more complex constraints, joint structures, and link geometries. The extent to which LLMs can contribute to design optimization depends on their spatial reasoning capabilities and the complexity of the problem setting. While this study primarily focused on demonstrating a new concept, future work will aim to develop a system that can be applied to more complex problem settings.

### v2

@@ Abstract
Index Terms:
Design Optimization, Multi-Objective Black-Box Optimization, Large Language Models
[…]
=-21pt
@@ I INTRODUCTION
As illustrated, various methods such as genetic algorithms, evolutionary strategies, exhaustive search, heuristic search, and GANs have been utilized for optimizing the body design of diverse robots. Among these, research using black-box optimization (BBO), particularly genetic algorithms, is prevalent due to its applicability to complex problem settings, including optimizations involving both continuous and discrete variables as well as multi-objective optimization [yang2000modular, xiao2016designopt, kawaharazuka2023autodesign, kawaharazuka2024slideopt, vanteddu2024cadurdf]. Although gradient-free black-box optimization is applicable to any problem, it has the drawback of requiring a substantial number of samples and time to generate high-quality solutions.
@@
To address this, we propose a method that combines black-box optimization with LLMs to enable more efficient output of design solutions. Our approach leverages the strength of black-box optimization for precise numerical adjustments and the capability of LLMs to produce broad design outputs. In parallel with the sampling process in conventional black-box optimization, sampling is also performed using an LLM, which is provided with prior knowledge, problem settings, and design examples. It should be noted that while the combination of LLMs and BBO has been shown to be effective for single-objective and general numerical problems [liu2024large], this study extends the approach to more complex robot design problems and further expands it from single-objective to multi-objective optimization. In this study, a multi-objective robot design optimization is conducted by adjusting the robot’s base position, joint configuration, and link lengths to minimize the end-effector position error and the required torque for given operation points. We demonstrate that our method enables more efficient exploration of design solutions compared to black-box optimization alone, while also discussing the method’s interesting characteristics and limitations.
The contributions of this study are as follows:
  • We propose a robot body design optimization method that combines large language models with multi-objective black-box optimization.
  • We demonstrate that the proposed method can explore design solutions more efficiently than black-box optimization alone.
@@ II Efficient Robot Design with Multi-Objective Black-Box Optimization and LLM-based Sampling
This study consists of the following four phases:
  • Definition of design parameters and constraints (Section II-A)
  • Computation of objective function values using inverse kinematics and torque evaluation (Section II-B)
  • Sampling with black-box optimization and a large language model at a fixed ratio (the first part of Section II-C and Section II-D)
  • Evaluation of optimization progress using hypervolume (the latter part of Section II-C)
@@ II-B Objective Functions
where $\bm{P}^{ik}_{i}$ and $\bm{\tau}^{ik}_{i}$ represent the end-effector position and gravity-compensation torque, respectively, when the robot solves the inverse kinematics (IK) to reach the target position $\bm{P}^{ref}_{i}$, assuming the robot’s initial posture has all joint angles set to zero. The inverse kinematics algorithm used in this study is based on the methods described in [chan1995ik, sugiura2007ik]. Joint angle limits are considered, whereas collision avoidance is not taken into account. The position error is defined as the minimum distance between the end effector and the target position achieved during 100 iterations of the inverse kinematics solver. Joint torques are computed by assuming each link to be a rigid body with uniform density. In other words, this evaluates whether the target end-effector position can be achieved with that body design, and how much joint torque is required in doing so. Of course, it is also possible to consider not only the target position but also the orientation, or to use a more complex evaluation function.
@@ II-C Multi-Objective Black-Box Optimization
In this study, we address a multi-objective optimization problem that aims to minimize both objective functions $E^{pos}$ and $E^{torque}$. This allows us to obtain various design solutions from the Pareto-optimal set. For the multi-objective optimization, we use the Tree-Structured Parzen Estimator (TPE) [bergstra2011tpe] implemented in the black-box optimization library Optuna [akiba2019optuna]. TPE supports multi-objective optimization and is notable for requiring a relatively small recommended sample size of fewer than 1000. Non-dominated Sorting Genetic Algorithm II (NSGA-II) [deb2002nsgaii] is also commonly used; however, it typically requires a larger number of samples and is less suitable than TPE for handling the integer and continuous parameters that frequently appear in design optimization problems.
Since comparing BBO algorithms is not the primary objective of this study, we do not include such comparisons in this work.
@@ III-A Experimental Setup
In this study, experiments are conducted for three sets of target operation points, Target-1, Target-2, and Target-3, as shown in Fig. 6. For each target operation point, we compare BBO, which uses only black-box optimization, and BBO/LLM+, which combines black-box optimization with a large language model. To evaluate the extent of LLM’s contribution to the optimization process, we vary $N_{step}$, which represents the number of black-box optimization steps performed between each LLM-based sampling. In this study, experiments are conducted with $N_{step}={5,10,50}$. Additionally, we perform a comparative evaluation of BBO/LLM-, which uses general Chain-of-Thought reasoning without specific parameter analysis instructions. The other parameters are set as follows: $D=4$, $N_{init}=10$, $N_{pareto}=5$, $N_{random}=5$, and $N_{total}=200$. GPT-4o was used as the LLM, with the temperature set to 0. Each experiment is repeated five times with different random seed values, and the mean and standard deviation of the results are reported.
@@ III-B Experimental Results
The experimental results are presented below. Fig. 7 shows a comparison of the methods BBO, BBO/LLM+ ($N_{step}=5$), BBO/LLM+ ($N_{step}=10$), and BBO/LLM+ ($N_{step}=50$) for Target-1, Target-2, and Target-3. The horizontal axis represents the optimization iterations, and the vertical axis represents the hypervolume $E_{hv}$. The figure illustrates how $E_{hv}$ changes with different values of $N_{step}$. A general trend observed is that BBO/LLM+ ($N_{step}=10$) consistently demonstrates the best results across all target operation points, confirming the effectiveness of LLM-based sampling. For Target-1, the improvement is particularly significant, while for Target-2 and Target-3, the impact is less pronounced, indicating that the effectiveness of LLM-based sampling depends on the target operation points. For $N_{step}=50$, only four LLM-based samplings occur within 200 sampling steps, making the trend of $E_{hv}$ similar to that of BBO. For $N_{step}=5$, although LLM-based sampling occurs more frequently and is effective in the early stages for Target-1 and Target-2, the results are inferior compared to $N_{step}=10$.
@@ III-D Target-2 Experiment
We describe the experimental results for Target-2 (Fig. 11). In the case of Target-2, LLM-based sampling also accounts for many of the Pareto-optimal solutions, confirming its effectiveness. Unlike Target-1, the joint structure for both Solution 1 and Solution 2 is the same, YRPRY-R-P-R. However, the second and fourth links in Solution 1 are longer compared to those in Solution 2. Due to this difference, Solution 1 achieves a lower $E^{pos}$, while Solution 2 achieves a lower $E^{torque}$. In Solution 2, the lengths of the first and second links are set to approximately 0.5 m, allowing the second link to match the height of the target operation points. The remaining joints are then used in a manner similar to a Selective Compliance Assembly Robot Arm-type (SCARA-type) structure, helping to minimize the required torque.
@@ IV Discussion
Finally, we discuss the limitation of the proposed method. In this study, we adopted a relatively straightforward and general problem setting. However, in real-world robot design, it is often necessary to consider more complex constraints, joint structures, and link geometries. The extent to which LLMs can contribute to design optimization depends on their spatial reasoning capabilities and the complexity of the problem setting. There remain several challenges with LLMs, including hallucinations, insufficient spatial reasoning capability, and high sensitivity to prompt variations. As these issues are addressed in future developments, the applicability of LLMs to design optimization is expected to further expand. While this study primarily focused on demonstrating a new concept, future work will aim to develop a system that can be applied to more complex problem settings. In addition, we believe that similar approaches can be extended to Neural Architecture Search (NAS) [zoph2017nas] and to the joint optimization of design and control, which we identify as promising directions for future research.